In England, a miniature electric bicycle has been constructed. The Ni-Cd batteries help the rider pedal for up to 3.00 h. If a charge of 1.51 x 10^4 C passes through the motor during those 3.00 h, what is the current?
1
Expert's answer
2011-02-07T06:55:22-0500
I = Q/t [C/sec]
3h = 3*60*60 = 10800 sec
I = 1.51x104 /10800 = 1.4 A.